# On Efficient Variants of Segment Anything Model: A Survey

### v2

@@ 3.1.2 Knowledge Distillation based Methods
Zhang et al. [159] made the first attempt to replace SAM’s heavy ViT encoder with the more efficient TinyViT [139], resulting in the integrated model named MobileSAM . As highlighted in [62], training SAM from scratch requires multiple days and 128 GPUs. MobileSAM attributes this complexity to the challenge of optimizing both the encoder and decoder simultaneously. To address this, they propose an encoder-only distillation strategy as illustrated in Fig. 8, which aims to transfer the visual representation capabilities of ViT-H to TinyViT. The loss function used is a simple Mean Square Error (MSE) between the output image embeddings of the two encoders. Further fine-tuning of the prompt encoder or mask decoder is optional and can lead to improved accuracy.
@@
Shu et al. [115] conducted an analysis of MobileSAM, identifying that encoder-only distillation can lead to significant performance degradation. To address this issue, they propose the Hard Mining Full-Stage Knowledge Distillation strategy for more effective distillation, as illustrated in Fig. 9. With a structure identical to MobileSAM, they developed a new SAM variant called TinySAM by training with this improved KD strategy. Specifically, the strategy supervises not only the image embeddings but also the output tokens and output masks, all using L1 Loss. To further enhance the distillation process, they introduced the Hard Mask Weighting strategy, which assigns larger weights to masks that are harder to predict, thereby improving learning efficiency. The factor $H$ is calculated as follows,
@@
Researchers from NVIDIA introduced a new SAM variant, NanoSAM [95], based on MobileSAM. It is designed to achieve real-time performance on NVIDIA Jetson Orin platforms using NVIDIA TensorRT. NanoSAM replaces the ViT-based encoder with a pure convolutional network, specifically ResNet18 [49], while retaining the other components from MobileSAM. NanoSAM is distilled from MobileSAM, and both models are retrained with TensorRT for optimized performance. The image encoder of MobileSAM is optimized with FP32 precision, whereas NanoSAM’s image encoder is built using FP16 precision for faster execution. Inference latency results on Jetson Orin Nano and Jetson AGX Orin demonstrate that NanoSAM is 5x faster than MobileSAM, with minimal accuracy loss.
Wang et al. [129] developed an efficient SAM variant, RepViT-SAM, using their newly proposed CNN-based backbone, RepViT [130], as the image encoder. The core idea behind RepViT is to integrate the effective design principles of efficient Vision Transformers (ViTs) into lightweight CNNs. These design principles are applied at three levels: block-level, macro-level, and micro-level. At the block level, RepViT separates the token mixer and channel mixer [155], reduces the expansion ratio [42], and increases the width of blocks. For the macro design, it incorporates early convolutions [141] as the input stem, deepens the down-sampling layers [83], employs a simpler classifier, and adjusts the block ratios across stages [102]. At the micro level, only 3x3 convolutions are used, and Squeeze-and-Excitation layers [54] are applied only in odd-numbered blocks. RepViT-SAM is trained using knowledge distillation, following the same pipeline as in [159], achieving a 10x increase in inference speed compared to MobileSAM.
@@
Zhang et al. [171] introduced EfficientViT-SAM, which utilizes EfficientViT [10] as the image encoder. The primary advantage of EfficientViT is its use of the ReLU linear attention mechanism, which facilitates global information interaction while enhancing hardware efficiency. By eliminating the hardware-unfriendly softmax operations and replacing them with ReLU, the attention calculation is reformulated as follows,
@@
Shen et al. [113] introduced FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [28], which extends the receptive field, with FlashAttention [25] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders, as shown in Fig. 11.
@@ 3.1.3 Quantization based Methods
Different from taking quantization in an encoder-only manner used for Q-TinySAM, Lv et al. [87] propose a novel framework that can directly take post-training quantization on SAM, namely PTQ4SAM. They start the work by first discovering the two challenges after traditional PTQ: 1) the existence of bimodal distribution which has negative effects on quantization quality and 2) the obvious discrepancy in the distribution of different attention mechanisms. Therefore, researchers bring up two strategies to solve them respectively: the Bimodal Integration (BIG) and the Adaptive Granularity Quantization (AGQ). For the Bimodal Integration, a sign factor $\gamma$ is introduced to convert the bimodal distribution into a normal distribution. For the Adaptive Granularity Quantization, the key is to adopt an adaptive parameter $\tau$ to adjust the base in the Log2 quantizer [92]. With the use of the AGQ strategy, the discrepancy between post-softmax distributions of different attention is expected to narrow down effectively. The PTQ4SAM is a plug-and-play SAM variant that can be easily deployed to downstream tasks.
@@ 3.2 Accelerating SegEvery Tasks
Shu et al. [115] observed that using a dense points grid (e.g., 32*32, 64*64) often generates a large number of redundant masks, which are later filtered out during post-processing, an operation that incurs significant time costs. In reality, only a few points within the grid are necessary to produce confident masks. To address this inefficiency, they proposed a hierarchical strategy for efficient sampling, which progressively selects optimal points for mask generation. This strategy involves two rounds of prompt generation. In the first round, a sparse grid is used, incorporating only a fraction (approximately 1/4) of the default points along each side. Masks are generated based on these points, and after filtering, only high-confidence masks are retained as final predictions. In the second round, a denser grid is applied, following the default configuration. However, points located in regions already covered by confident masks are excluded, significantly reducing the number of points. The predicted results from both rounds are then fused to produce the final output. The pipeline for this hierarchical strategy is illustrated in Fig. 13. By employing this two-round approach, the sampling process becomes both more time-efficient and fine-grained, leading to a notable acceleration in the SegEvery task with minimal performance degradation.
@@ 3.3 Future Research Directions
Exploring Advanced Architectures. While current SAM variants have demonstrated efficiency gains through the adoption of efficient architectures and model compression techniques, there is substantial potential for further improvement. The exploration of Transformer-alternative models, such as Mamba [43], RetNet [117], KAN [84], and TTT [118], presents an exciting opportunity to design more lightweight and efficient structures. These models may offer advantages in computational efficiency without sacrificing segmentation accuracy. In addition to alternative models, refining the efficiency of the attention mechanism across both the image encoder and mask decoder remains critical. Methods such as linear attention, low-rank factorization, or hybrid architectures combining convolutional and attention-based designs should be further investigated. Addressing bottlenecks in both computation and memory usage will enhance SAM’s deployment across diverse hardware environments.
@@
Exploiting Sparsity and Acceleration Techniques. The inherent sparsity observed in deep neural networks, where only a subset of parameters significantly contributes to model output, offers a promising avenue for improving SAM’s efficiency. Techniques like pruning, quantization, and structured sparsity could further reduce SAM’s computational demands. While initial efforts in sparsification, such as those in [20], have shown success, future research could focus on understanding the distribution and dynamics of sparsity in SAM’s architecture. This includes investigating the optimal layers or components of SAM that can be pruned or sparsified without impacting performance. Additionally, techniques such as sparse attention mechanisms, dynamic pruning during inference, and low-precision training should be explored to balance accuracy and efficiency, especially for large-scale deployments. By combining these with advanced knowledge distillation techniques, more compact, efficient variants of SAM could be realized.
@@
Hardware-Specific Optimizations. Optimizing SAM for specific hardware platforms, including GPUs, TPUs, specialized AI accelerators (e.g., NVIDIA’s TensorRT or Google’s Edge TPU), and edge devices, can unlock significant gains in performance and efficiency. Hardware-aware model optimization techniques, such as operator fusion, quantization-aware training, and custom CUDA kernels, can maximize throughput and reduce latency when deploying SAM on modern hardware platforms. In the context of edge devices, which face extreme constraints in storage, computational power, and energy supply, these optimizations are crucial for real-time applications, such as segmentation on UAVs or IoT devices. Future research could explore hierarchical cloud-edge architectures to offload computationally expensive tasks to the cloud, while running lightweight models locally on edge devices. Additionally, leveraging specialized AI hardware like Field-Programmable Gate Arrays (FPGAs) or using techniques such as hardware-aware neural architecture search (NAS) and mixed-precision quantization could further optimize SAM for low-latency, resource-constrained environments, ensuring that the model operates effectively across diverse hardware platforms.
@@
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [105] for video segmentation and MM-SAM [140] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation and cross-modal feature sharing could significantly reduce computational costs. A recent work, named Surgical SAM 2 [77], has proposed the efficient frame pruning strategy which only retains the most informative frames in memory bank to achieve reduction in memory usage and acceleration in inference, providing a promising idea for efficient segmentation in videos. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention, efficient memory mechanisms for temporal data, and early-fusion techniques to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.

### v3

@@ 3.1.2 Knowledge Distillation based Methods
Zhang et al. [43] made the first attempt to replace SAM’s heavy ViT encoder with the more efficient TinyViT [114], resulting in the integrated model named MobileSAM . As highlighted in [19], training SAM from scratch requires multiple days and 128 GPUs. MobileSAM attributes this complexity to the challenge of optimizing both the encoder and decoder simultaneously. To address this, they propose an encoder-only distillation strategy as illustrated in Fig. 8, which aims to transfer the visual representation capabilities of ViT-H to TinyViT. The loss function used is a simple Mean Square Error (MSE) between the output image embeddings of the two encoders. Further fine-tuning of the prompt encoder or mask decoder is optional and can lead to improved accuracy. The image encoder of MobileSAM has been further applied to HQ-SAM [104], a variant significantly boosting SAM’s mask quality, for mobile efficiency. The resulted model, namely light HQ-SAM, exceeds MobileSAM in performance with negligible gap in inference speed, proving it an excellent SAM variant trading off well between accuracy and efficiency.
@@
Not long after MobileSAM’s publishing, Shu et al. [44] conducted an analysis on it, identifying that encoder-only distillation can lead to significant performance degradation. To address this issue, they propose the Hard Mining Full-Stage Knowledge Distillation strategy for more effective distillation, as illustrated in Fig. 9. With a structure identical to MobileSAM, they developed a new SAM variant called TinySAM by training with this improved KD strategy. Specifically, the strategy supervises not only the image embeddings but also the output tokens and output masks, all using L1 Loss. To further enhance the distillation process, they introduced the Hard Mask Weighting strategy, which assigns larger weights to masks that are harder to predict, thereby improving learning efficiency. The factor $H$ is calculated as follows,
@@
Researchers from NVIDIA introduced a new SAM variant, NanoSAM [138], based on MobileSAM. It is designed to achieve real-time performance on NVIDIA Jetson Orin platforms using NVIDIA TensorRT. NanoSAM replaces the ViT-based encoder with a pure convolutional network, specifically ResNet18 [112], while retaining the other components from MobileSAM. NanoSAM is distilled from MobileSAM, and both models are retrained with TensorRT for optimized performance. The image encoder of MobileSAM is optimized with FP32 precision, whereas NanoSAM’s image encoder is built using FP16 precision for faster execution. Inference latency results on Jetson Orin Nano and Jetson AGX Orin demonstrate that NanoSAM is 5x faster than MobileSAM, with minimal accuracy loss. Very recently, a similar work, which builds up an ultra light-weight variant based on the U-Net architecture, named PicoSAM2 [151], has successfully deployed SAM on Sony IMX500 (a camera sensor with 8MB memory limit.
Due to MobileSAM with TinyViT as backbone still encounters challenges on resource-constrained devices, Wang et al. [113] developed an novel efficient SAM variant, RepViT-SAM, using their newly proposed CNN-based backbone, RepViT [45], as the image encoder. The core idea behind RepViT is to integrate the effective design principles of efficient Vision Transformers (ViTs) into lightweight CNNs. These design principles are applied at three levels: block-level, macro-level, and micro-level. At the block level, RepViT separates the token mixer and channel mixer [183], reduces the expansion ratio [184], and increases the width of blocks. For the macro design, it incorporates early convolutions [185] as the input stem, deepens the down-sampling layers [186], employs a simpler classifier, and adjusts the block ratios across stages [187]. At the micro level, only 3x3 convolutions are used, and Squeeze-and-Excitation layers [188] are applied only in odd-numbered blocks. RepViT-SAM is trained by following the same distillation pipeline in [43], claiming to achieve a 10x increase in inference speed compared to MobileSAM.
@@
To further improve the efficiency of backbone while maintaining strong capability of feature extraction, Zhang et al. [122] introduced EfficientViT [116] as image encoder, building up the powerful variant, EfficientViT-SAM. The primary advantage of EfficientViT is its use of the ReLU linear attention mechanism, which facilitates global information interaction while enhancing hardware efficiency. By eliminating the hardware-unfriendly softmax operations and replacing them with ReLU, the attention calculation is reformulated as follows,
@@
Following the similar idea of applying high-efficiency attention mechanisms, Shen et al. [128] proposed FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [189], which extends the receptive field, with FlashAttention [127] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders, as shown in Fig. 11.
@@ 3.1.3 Quantization based Methods
Instead of taking quantization in an encoder-only manner used for Q-TinySAM, Lv et al. [47] propose a novel framework that can directly take post-training quantization on SAM, namely PTQ4SAM. They start the work by first discovering the two challenges after traditional PTQ: 1) the existence of bimodal distribution which has negative effects on quantization quality and 2) the obvious discrepancy in the distribution of different attention mechanisms. Therefore, researchers bring up two strategies to solve them respectively: the Bimodal Integration (BIG) and the Adaptive Granularity Quantization (AGQ). For the Bimodal Integration, a sign factor $\gamma$ is introduced to convert the bimodal distribution into a normal distribution. For the Adaptive Granularity Quantization, the key is to adopt an adaptive parameter $\tau$ to adjust the base in the Log2 quantizer [190]. With the use of the AGQ strategy, the discrepancy between post-softmax distributions of different attention is expected to narrow down effectively. The PTQ4SAM is a plug-and-play SAM variant that can be easily deployed to downstream tasks.
Concurrent with PTQ4SAM, a counterpart named PQ-SAM [152], further exploits the post-training quantization method customized for SAM’s activations, achieving the first usable 4-bit PTQ for SAM. The core idea lies in transferring SAM’s highly asymmetric activation distribution with extreme outliers into quantization-friendly distribution, for which the Grouped Activation Distribution Transformation (GAPT) is proposed. GAPT firstly adopts tensor-wise outlier truncating to eliminate outliers in activation values, followed by the channel-wise outlier grouping to classify channels of truncated tensors into clusters based on their range. Pairs of learnable shifting and scaling sizes are subsequently allocated to each channel group, reshaping the distribution as follows:
$\tilde{X}(c)=(X(c)-{v_{n}})\oslash{s_{n}}$ (9)
[…]
@@ 3.1.4 Pruning based Methods
Unlike Chen et al. [48] adopting the pruning and post-alignment strategy to lighten SAM in a straightforward manner, Abebe et al. [153] incorporates pruning methods with the one-shot Neural Architecture Search [191] (NAS, an effective technique of automating the design of efficient neural networks), building up a weight-sharing supernetwork of SAM, named SuperSAM. The overall architecture of SuperSAM is illustrated as Fig. 13. SuperSAM works as a search space of SAM-like subnetworks with size 30-70% smaller than original SAM-B, while some of them still maintaining considerable performance are considered as efficient variants of SAM. The supernetwork is trained by sampling and training subnetworks with flexible image encoder, in which the structured pruning is applied to remove insignificant Transformer layers, with further weight reordering and slicing of MLP in remaining layers. Once the SuperSAM is built up, techniques such as OpenTuner [192] can be deployed to search for the architecture meeting both efficiency and accuracy requirement. The emerge of SuperSAM might further broaden the domain of investigating efficient SAM variants.
Figure 13: The architecture of SuperSAM [153] consists of a flexible image encoder and SAM-like prompt-guided mask decoder. Lightweight SAM variants can be built up by sampling appropriate image encoder meeting predefined efficiency and accuracy requirements.
@@ 3.1.5 Code Refactorization
3.1.6 Discussion
In this part, we discuss the trade-offs among different architectures and methodologies, and suggest best practices for applying efficient SAM variants across various conditions. For SAM variants, the choice of backbone largely determines the model size, computational complexity, and segmentation performance. These backbones can be broadly classified into three groups: CNN‐based, ViT‐based, and attention‐modified architectures. CNN‐based backbones leverage small and parameter‐shared kernels for feature extraction, benefiting with fewer parameters and well-optimized hardware deployment, leading to efficient training and inference; however, their inherent locality limits the capture of global context and, consequently the segmentation accuracy. In contrast, ViT‐based structures overcome this limitation by employing attention mechanisms that enhance the context of feature representations and boost overall segmentation accuracy. This advantage comes at the cost of heavy training cost and quadratic computational complexity, rendering them computationally expensive and less efficient for reasoning. Attention‐modified backbones address the inefficiencies of traditional attention by either introducing more computationally efficient operations or by replacing the attention mechanism with Transformer‐alternatives, thereby reducing complexity while maintaining acceptable accuracy. However, efficient attention mechanisms may still suffer from challenges by hyperparameter sensitivity and implementation complexity.
Combining different architectural components within SAM variants might further overcome the above deficiencies by complimentarily taking advantages from each architecture, leading to better balance between efficiency and accuracy. Intuitively, hybrid designs that integrate CNN-based structures with Transformer-based modules or Transformer-alternative models are feasible solutions. There have been attempts in recent SAM variants, such as Lite-SAM [124], which employs a CNN-Transformer hybrid encoder LiteViT, and RWKV-SAM [136], which integrates CNNs with the RWKV model. These variants with hybrid architectures have demonstrated significant improvements in efficiency while maintaining comparable segmentation performance to vanilla SAM. Despite their promise, the mismatch between heterogeneous layers, which may lead to low quality in feature representations and high difficulty in model optimization, is a non-negligible challenge for building up effective hybrid models, and as a result, more advanced designs about module combination are needed in the future research.
In addition to backbone selection, additional compression techniques such as quantization and pruning can further reduce model size and memory usage during inference. However, these methods may not always deliver the expected runtime speedup and can sometimes lead to a significant degradation in accuracy and generalization. The training procedure itself also has a profound impact on performance. Training from scratch can yield comprehensive, high-quality models but requires substantial training resources. Knowledge distillation offers a more efficient alternative using less data. Yet, distilled models may fail to fully inherit the teacher’s capabilities, resulting in performance inferior to models trained from scratch. Moreover, combinations of these methods may yield models with varied accuracy–efficiency profiles. For example, pruning combined with post-quantization can compress a model to an extremely lightweight variant by reducing both parameter size and model complexity; subsequent fine-tuning (either through full training or distillation) can then help recover lost accuracy, thereby achieving a more balanced design.
Based on the analysis above and the experimental results in Sections 4.2 and 4.3, we suggest several best practices for deploying efficient SAM variants across diverse hardware platforms. In GPU environments, where computational resources are relatively abundant, thus throughput are the primary concerns. Models like EfficientViT-SAM offers an excellent balance between speed and segmentation accuracy. Its streamlined backbone featuring linear ReLU attention ensures the model capacity with reasonable memory cost, benefiting from the two-stage training procedure (initial distillation followed by full training) further enhances segmentation performance. Conversely, in CPU settings where parallel matrix computations are limited and models with high computational complexity can lead to prohibitive inference latencies, efficiency becomes critical. Architectures like NanoSAM with lightweight CNN-based backbones can theoretically minimize computational complexity. Our runtime evaluations on CPUs indicate that EfficientViT-SAM-L0 is the most effective variant, combining rapid inference with sufficient segmentation accuracy. For edge devices, which typically operate under stringent resource constraints and demand real-time performance, NanoSAM is the preferred option for minimizing memory usage and reducing latency. However, if the available hardware supports a slightly heavier model without compromising real-time requirements, EdgeSAM also emerges as viable alternatives that provide a more balanced combination of speed and accuracy.
@@ 3.2 Accelerating SegEvery Tasks
Meanwhile, Shu et al. [44] also observed that using a dense points grid (e.g., 32*32, 64*64) often generates a large number of redundant masks, which are later filtered out during post-processing, an operation that incurs significant time costs. In reality, only a few points within the grid are necessary to produce confident masks. To address this inefficiency, they proposed a hierarchical strategy for efficient sampling, which progressively selects optimal points for mask generation. This strategy involves two rounds of prompt generation. In the first round, a sparse grid is used, incorporating only a fraction (approximately 1/4) of the default points along each side. Masks are generated based on these points, and after filtering, only high-confidence masks are retained as final predictions. In the second round, a denser grid is applied, following the default configuration. However, points located in regions already covered by confident masks are excluded, significantly reducing the number of points. The predicted results from both rounds are then fused to produce the final output. The pipeline for this hierarchical strategy is illustrated in Fig. 14. By employing this two-round approach, the sampling process becomes both more time-efficient and fine-grained, leading to a notable acceleration in the SegEvery task with minimal performance degradation.
@@
Similar to Lite-SAM’s idea of automatic prompt generation, the recent variant AoP-SAM proposed by Chen et al. [156], employs an efficient Prompt Predictor module to identify the optimal regions for prompt candidates, followed by an Adaptive Sampling and Filtering (ASF) module to generate prompts in a coarse-to-fine manner. The architecture of AoP-SAM is illustrated in Fig. 16. The Prompt Predictor consists of two encoder, processing original images and embedding from SAM’s image encoder respectively, and one decoder, incorporating the spatial and context information from encoders to generate the prompt confidence map. The function of ASF module is to extract essential prompts from prompt confidence map, which firstly samples the local maxima as coarse prompts, and secondly filters by the IoU scores of previously generated referenced masks and elimination scores of prompts, ensuring only crucial prompts will be retained for later mask generation. The elimination score is calculated by referenced mask and image embedding, indicating the possibility of being redundant. The experiments coudcted on three image segmentation datasets, have demonstrated AoP-SAM’s highly efficiency for SegEvery task, surpassing both SAM’s dense grid sampling strategy and MobileSAMv2’s [155] object-aware sampling strategy.
Figure 16: The architecture of AoP-SAM [156]. It adopts a Prompt Predictor to predict prompt confidence map and an Adaptive Sampling and Filtering (ASF) module to generate final prompts.
@@ 3.3 Future Research Directions
Exploring Advanced Architectures. While current SAM variants have demonstrated efficiency gains through the adoption of efficient architectures and model compression techniques, there is substantial potential for further improvement. The exploration of Transformer-alternative models, such as Mamba [133], RetNet [193], KAN [194], and TTT [195], presents an exciting opportunity to design more lightweight and efficient structures. With the success of incorporating RWKV into SAM [136], leveraging Transformer-alternative models as backbone has shown great potential to build up efficient SAM variants, which may offer advantages in computational efficiency without sacrificing segmentation accuracy. In addition to alternative models, refining attention modules’ efficiency remains critical for further development of accelerating SAM, inspired by the preliminary exploration of FastSAM3D [128]. Methods such as linear attention, low-rank factorization, or hybrid architectures combining convolutional and attention-based designs, which have been widely applied to Transformer-based models [196, 163, 197], should be further investigated for efficient SAM. Addressing bottlenecks in both computation and memory usage will enhance SAM’s deployment across diverse hardware environments.
@@
Exploiting Sparsity and Acceleration Techniques. The inherent sparsity observed in deep neural networks, where only a subset of parameters significantly contributes to model output, offers a promising avenue for improving SAM’s efficiency. The discovery in very recent work, SuperSAM [153], showing that the subnetwork with 7 layers in image encoder pruned (12 layers in total) can still maintain over 60% performance compared to vanilla SAM, further provides proofs to SAM’s sparsity, which poses a promising research direction for further uncovering the fine-grained sparsity in SAM’s architecture. Techniques like pruning, quantization, and structured sparsity [154] could further reduce SAM’s computational demands. While several initial efforts in sparsification, such as those leveraged in SlimSAM [48] and PQ-SAM[152], have shown success, future research could dive deeper into understanding the distribution and dynamics of sparsity in SAM. This includes investigating the optimal layers or components of SAM that can be pruned or sparsified without impacting performance [153]. Additionally, techniques such as sparse attention mechanisms, dynamic pruning during inference, and low-precision training, which have been deployed on Transformer-based architectures [198, 199, 200], can be further explored to balance accuracy and efficiency, especially for large-scale deployments. By combining these with advanced knowledge distillation techniques, more compact, efficient variants of SAM could be realized.
@@
Hardware-Specific Optimizations. Optimizing SAM for specific hardware platforms, including GPUs, TPUs, specialized AI accelerators (e.g., NVIDIA’s TensorRT or Google’s Edge TPU), and edge devices, can unlock significant gains in performance and efficiency. Equipping SAM with hardware-aware model optimization techniques, for instance, operator fusion, quantization-aware training, and custom CUDA kernels, some of which have been applied in SAMfast [154] by PyTorch extensions, requires further exploration in the future to maximize throughput and reduce latency when deploying SAM on modern hardware platforms. In the context of edge devices, which face extreme constraints in storage, computational power, and energy supply, these optimizations are crucial for real-time applications, such as segmentation on UAVs or IoT devices. Future research could explore hierarchical cloud-edge architectures to offload computationally expensive tasks to the cloud, while running lightweight models locally on edge devices. This idea has already been investigated in LLMs for robot control [201]. Additionally, leveraging specialized AI hardware like Field-Programmable Gate Arrays (FPGAs) [202] or using techniques such as hardware-aware neural architecture search (NAS) and mixed-precision quantization [203] might further optimize SAM for low-latency, resource-constrained environments, ensuring that the model operates effectively across diverse hardware platforms. Among them, NAS is now successfully applied to SAM, building up the first supernetwork of SAM [153].
@@
Universal Segmentation among Multiple Domains. The development of efficient SAM variants has been witnessed to achieve a great success, achieving significant improvement in runtime efficiency while demonstrating reasonable performance in image-level segmentation tasks. However, most of these works only conduct accuracy evaluation on COCO and LVIS datasets, neglecting to assess efficient variants’ universal segmentation capability across various domains. To mitigate this gap, in our survey, we further evaluate models on two diverse, multi-domain benchmarks, Segmentation in the Wild (SGinW) and Unidentified Video Objects (UVO), to strengthen the evaluation towards SAM variants’ generalization performance in varied scenarios. Results illustrated in Tab. 10 and Tab. 11 have shown that SAM variants may demonstrate disastrous performance when handling uncommon objects, for instance, several models falling short to segment Poles and Rail in SGinW benchmark, indicating efficient SAM variants’ deficiency in rarely seen or unseen targets. However, our experiments still do not reach the perfection, for not covering assessment of these variants’ performance on some domain-specific tasks, such as segmentation in medical images or aerial images, which may further reveal their weakness in universal generalization. Future work might focus on building up an efficient, almighty variant of SAM, boosting both the runtime efficiency and overall performance across a broader range of downstream tasks.
@@
Efficient Segmentation for Video and Multi-Modality Data. Video and multi-modality tasks, which deal with complex, dynamic environments, are rapidly gaining relevance across numerous real-world applications. While some initial efforts, such as SAM 2 [34] for video segmentation and MM-SAM [209] for multi-modality tasks, have extended SAM’s applicability, efficiency remains a pressing issue. Video data contains temporal redundancy, while multimodal data often exhibits correlations between modalities. Leveraging these inherent redundancies through techniques like temporal aggregation [210] and cross-modal feature sharing [211] could significantly reduce computational costs. Future work could focus on optimizing SAM’s runtime complexity by utilizing spatiotemporal attention [212], efficient memory mechanisms (which are now effective solutions for accelerating SAM 2 [213, 214]) for temporal data, and early-fusion techniques [215] to reduce the number of modality-specific computations. Developing frameworks that adapt dynamically to different levels of redundancy across frames or modalities can further drive computational efficiency in real-world applications.
Efficient Variants of SAM 2. Segment Anything Model 2, successor of Segment Anything Model, which not only achieves higher accuracy and efficiency on image segmentation task but also extends its powerful capability into video segmentation tasks, has been increasingly gaining attention from industry and research community. However, similar to SAM’s dilemma, SAM 2 also suffers the inefficiency issue due to the large hierarchical image encoder, hindering its application in various resource-constrained, real-world scenarios. Moreover, another main bottleneck of SAM 2 lies in the newly introduced memory mechanism, more specifically, the memory attention module which attends feature map to memories, as memory features holds the same spatial size to image feature, resulting in long memory token sequence and heavy computational cost in cross-attention. Several works of efficient SAM 2 [216, 213, 214] has emerged focusing on solving these two bottlenecks. Surgical SAM 2 [216] has proposed the efficient frame pruning strategy which only retains the most informative frames in memory bank to achieve reduction in memory usage and acceleration in inference, providing a promising idea for efficient segmentation in videos. The subsequent variant, named EfficientTAM [213], has built up an efficient memory module, which reconstructs memories by firstly adopting average pooling in local window and secondly restoring the spatial size through repeating, leading to less computation in memory attention. Furthermore, EfficientTAM replaces the image encoder with SAMI [109] pre-trained vanilla ViT, achieving higher efficiency and maintaining competing performance. Recently, a mobile-friendly SAM 2 variant, named EdgeTAM [214], further speeds up the memory attention by leveraging a global Perciever [217] and a novel 2D Spatial Perceiver to effectively compress memories. Moreover, EdgeTAM opts for RepViT as backbone and retrains through a two-stage distillation pipeline with SAM 2 as teacher model, pursuing for efficiency and accuracy improvements. Quantization techniques have also been preliminarily employed in SAM 2, for instance, the newly proposed variant, Q-SAM2 [218], which introduces calibration in SAM 2’s image encoder and subsequently implements quantization-aware training pipeline, effectively compressing model into low-bit configurations. Compared to SAM, there are much less works on efficient variants for SAM 2 up till now, posing it a challenging task. Techniques and strategies proved to be effective in accelerating SAM, have not been fully investigated on SAM 2, which might be the potential direction for future research.